Let X be a normed linear space and Y a closed subspace of XΒ 

with Y β‰  X ,if 0 < π‘Ÿ < 1 prove that there existΒ 

𝑋r element of X such that ||X|| = 1And π‘Ÿ < 𝑑 π‘‹π‘Ÿ

, π‘Œ ≀ 1


1
Expert's answer
2022-01-31T16:07:06-0500

First of all, let us remark that

"d(x_r, Y) = \inf_{y\in Y} d(x_r, y) \leq ||x_r||", as "0\in Y" and "d(0,x_r)=||x_r||". Therefore, the right-hand side of the inequality is automatically satisfied if "||x_r||=1".

Secondly, let us remark the following facts :

  1. For a constant "\lambda\in \mathbb{R}" and any "x\in X" we have "d(\lambda x, Y)=|\lambda|d(x,Y)"
  2. For a vector "y\in Y" and any "x\in X" we have "d(x-y, Y)= d(x,Y)"


We know that "x\mapsto d(x, Y)" is not a function that is constantly zero (as "Y" is closed). Therefore there exists "x_0" such that "d(x_0, Y) = a>0". By "\inf" property for any "\epsilon>0", there exists "y_\epsilon \in Y" such that "a\leq||x_0-y_\epsilon||\leq a+\epsilon". Now let us take the vector "x_\epsilon = \frac{x_0-y_\epsilon}{||x_0-y_\epsilon||}", by using the properties 1 and 2 we have

"1-\frac{\epsilon}{a+\epsilon}\leq d(x_\epsilon, Y) \leq 1"

Let us fix "0<r<1". Now by taking "\epsilon>0" small enough, taking into the account that "a>0" (and so "\lim_{\epsilon\to 0} \frac{\epsilon}{a+\epsilon} = 0") , we have "1-\frac{\epsilon}{a+\epsilon}>r" and thus "r<d(x_\varepsilon,Y)\leq 1" with "||x_\epsilon||=1"